# Interpretable Retinal Disease Prediction Using Biology-Informed Heterogeneous Graph Representations

### v1

@@ III-C1 Attribution Method
Instead of using a constant baseline sample, such as a zero-vector, we propose a dynamic baseline [45] that considers the unique characteristics of retinal vasculature, generating more fine-grained explanations. In contrast to natural images, the retinal vasculature adheres to a fixed structure that can be represented in the explanation process. Previous studies indicate variations in capillary density with increasing distance from the FAZ and across different retinal sectors [46]. This implies that distinct characteristics should be considered normal for intercapillary areas and vessels in different areas of the retinal vasculature. Therefore, we select the k-nearest same-type neighbor nodes from the training set and use their feature-wise median as the dynamic baseline vector $x^{bl}$ for each node in a graph. Formally, the baseline vector $x_{v}^{bl}$ for a node $v$ is defined as
$x_{v}^{bl}=\text{{median}}\left(\left\{x_{u}\,\middle|\,u\in\mathcal{N}_{l}(v)\right\}\right).$ (5)
@@ III-C2 Localization Method
For clinical utility, it is critical that graph nodes that have been identified as important (e.g., individual vessel segments) can be visualized in the corresponding OCTA image. These visualizations can be used to verify whether the extracted retinal characteristics align with the underlying image or if they arise from representation errors, such as image artifacts. We store the corresponding segmentation map for each node in our graph representation to achieve such a localization.
III-C3 Explanation Generation
@@
To generate our explanations, we first identify critical nodes using our attribution method by summing over the integrated gradients for the feature vector. The identified critical nodes are then spatially highlighted using the described localization method. Lastly, we identify the decisive characteristics of each critical node using the feature-wise integrated gradients. Examples of the generated spatial and feature-wise explanations are described in Section IV and visualized in Figure 5.
III-D Datasets
III-D1 Proprietary Clinical Dataset
@@
Our OCTA dataset contains 1268 OCTA images of the DVC. For some patients, images of both the right and left eye are contained in the dataset. Beyond that, no patients are included more than once. DR grades are assigned on a patient level with the stages of healthy (847 samples), diabetes mellitus (DM) without DR (133 samples), early non-proliferate DR (early NPDR, 134 samples), late NPDR (62 samples), and PDR (92 samples). The images are 3 mm $\times$ 3 mm, with a resolution of 304$\times$304 pixels.
@@
We perform a stratified data split into six parts for our experiments. As sometimes both eyes of a single patient are imaged, we strictly keep both samples in the same split to avoid information leakage. One of the splits is kept as a test set, while the other five splits are used for cross-validation. In each split, healthy and DM are pooled into a single healthy class, and early NPDR and late NPDR are pooled into a single NPDR class. As is typical in a real-world setting, our datasets are imbalanced towards healthy subjects. For these reasons, we evaluate and select our models based on balanced accuracy.
@@
TABLE II: Classification performance for the staging task into healthy, NPDR and PDR. The data representation describes the information used as input to the subsequent classifier. Results are obtained on a test set that was separated before cross-validation. Standard deviations are calculated based on the results from the 5-fold cross-validation. Our model performs best in the most important metric, i.e., balanced accuracy. We report p-values for paired t-tests with the heterogeneous graph model in parentheses after the standard deviations.
@@
Onset ROC
AUC
Trad. ML Tabular Biomarkers RF
@@
Transformer Image RETFound Fund.
@@
Trad. ML Node Embeddings RF
@@
III-D2 Public Dataset
We use the OCTA-500 [47] dataset as an external validation dataset for our method. For that purpose, we use the 200 3 mm $\times$ 3 mm images with a resolution of 304$\times$304 pixels and remove samples that contain labels other than DR or heatlhy. This results in a total of 189 images, of which we use the ILM-OPL projection maps. Notably, these projections contain SVC and DVC vessels, resulting in a slight domain shift compared to our proprietary dataset. Of the 189 images, 160 are healthy samples and 29 samples are in the DR class. We apply each algorithmic class’s best-performing checkpoints to classify the OCTA-500 samples. Predictions of NPDR and PDR are pooled into a single DR class for comparison with the ground truth labels.
IV DR Staging Experiments
The DR stage label is assigned based on the International Clinical Disease Severity Scale [48] relying on color fundus photography. We want to stress that this feature-based grading approach does not always align with the microvasculature changes visible in OCTA images. Recognizable changes in the microvasculature can occur e.g. before the presence of microaneurisms that result in the NPDR stage. Therefore, the achieved classification scores deviate strongly from perfect classification in terms of the International Clinical Disease Severity Scale.
IV-A Baseline Methods
We compare our approach to the most prominent and well-performing approaches for DR staging in the literature. These methods include manual biomarker extraction, and DL approaches such as CNNs directly operating on the information-rich raw OCTA images. Furthermore, we compare our method to a recently published foundational model that has not previously been employed for this task on OCTA images. Finally, we benchmark against traditional ML methods on graphs that employ global feature aggregation followed by classic ML classifiers without using graph convolutions.
TABLE III: Classification performance for the binary classification into healthy versus DR. The results are obtained on the external OCTA-500 without further fine-tuning the models previously trained on the other reported dataset. The best-performing checkpoint from the 5-fold cross-validation on the proprietary dataset is used for evaluation on the OCTA-500 dataset.
@@ IV-A Baseline Methods
Trad. ML Tabular Biom. RF 0.821 0.963 0.760 0.960 0.872
@@
IV-A1 Manual Biomarker Extraction
We extract traditional biomarkers from the image segmentations as a means of comparison. We use RF and SVM as classifiers for the DR stage prediction due to their excellent performance in a wide range of applications for tabular data. The biomarkers that we extract and use as classification features are namely:
  1. 1.
    FAZ biomarkers: area, maximal diameter, mean diameter, acircularity index, STD4, and NR300. The reader is referred to [49] for detailed descriptions of the last three biomarkers.
  2. 2.
    Vessel and intercapillary area biomarkers: vessel density, vessel perimeter, and fractal dimension as vessel complexity measure.
IV-A2 Vision Models
We compare against well-established CNN architectures previously used for the task at hand [50]. We implemented EfficientNet [18], ResNet [16], and VGG [17] models in combination with an optimized data augmentation pipeline. We use all models with weights pre-trained on ImageNet. Beyond these networks, we use the more recent ConvNeXt [51], Efficient Net V2 [52], and SwinTransfomer [53] architectures. We use cross-entropy loss for CNNs and all other neural network architectures.
IV-A3 RETFound Foundational Model
The RETFound model [19] is built on a vision transformer architecture. To initialize this complex architecture’s weights, the model is pre-trained in a self-supervised fashion on more than 900,000 color fundus photography images before being fine-tuned on our training set. We decided to use the color fundus pre-trained model instead of the OCT fine-tuned models due to better performance. This is likely caused by the larger domain shift from en-face OCTA to OCT than to color fundus photography.
IV-A4 Traditional ML on Node Embeddings
IV-B Results
IV-B1 Comparison of Graph Learning with Baselines
Our biology-informed heterogeneous graph representation, combined with a shallow GNN model, outperforms all baseline algorithms with a balanced accuracy of 69.7% (see Table II). It also obtains the highest ROC-AUC, surpassing all alternative approaches. Among vision models, the highest ROC-AUC scores were achieved by the ConvNeXt-s and VGG11 architectures. In comparison, the finetuned RETFound model only achieves a performance of 63.8%, which can be attributed to the large domain gap between the pretraining (color fundus images) and fine-tuning data (OCTA). Although the model learns about retinal characteristics during pretraining, it cannot transfer this knowledge to OCTA images, which capture fine-grained capillary structures and exhibit different colors and contrasts. Additionally, the complex underlying vision transformer architecture has a high data demand during training, which the available fine-tuning dataset does not satisfy. This observation is further supported by the comparably low performance of the SwinTransformer model. In contrast, our graph representation enables using a lightweight architecture (see Table I) that achieves peak performance even with limited data. Furthermore, the relative performance gap between the best CNN model and the best biomarker-based approach of $\sim$ 3% supports earlier findings that biomarker-based approaches cannot extract all relevant information necessary to achieve optimal performance.
Figure 5: Explanations for the decision-making of our proposed disease staging utilizing graph representations and graph classification. Rows a) and b) represent two separate samples that were correctly classified as NPDR and PDR respectively. The left-most images represent the original OCTA images. The pictures to the right spatially highlight the vessels (red) and intercapillary areas (blue) that are most important to the classification. The FAZ (green) is highlighted if it is among the most important nodes in the graph. In the neighboring barplot, the most important properties of the respective entities are displayed according to their relative importance. The right-most information about the deviation from the mean indicated if an increased or decreased value for the characteristics is causing the model to predict the stated disease stage.
IV-B2 Comparison of Random Forest on Node Embeddings to Graph Learning
Notably, a simple random forest classifier that operates on all aggregated node embeddings yields excellent results and even surpasses some CNNs and all traditional biomarker-based baselines. The ROC-AUC for the RF on the node embeddings (0.828) is significantly higher compared to the RF on the biomarkers (0.796), with a p-value $<0.01$. This observation shows that the nodes’ feature embeddings, consisting of geometric descriptors and intensity statistics, contain valuable information for downstream tasks. Furthermore, it supports the finding that tree-based methods outperform neural networks on tabular data tasks in various settings [54]. However, our method performs superiorly because the graph learning framework exploits the graph representation’s message-passing capabilities.
IV-B3 Comparison to SVC and Segmentation Maps
@@
A performance comparison for using the SVC instead of the DVC using the same CNN architecture shows similar staging performance for using either the DVC or the SVC layer. Moreover, we experimented with directly using segmentation masks as input to the CNN classifier. Interestingly, this results in comparable performance to raw images and opens an interesting avenue for enhancing explainability in classification with CNN models.
IV-B4 Generalization to OCTA-500
@@
We use the best-performing checkpoint from the 5-fold cross-validation on the proprietary dataset for evaluation on the OCTA-500 dataset. Our proposed graph learning method performs best in this 0-shot evaluation on the OCTA-500 dataset (see Table III). This highlights the robustness of our approach to minor domain changes induced by different hardware (OCTA scanners) and preprocessing algorithms (e.g., layer segmentation and intensity clipping). The performance between the CNN models and the RETFound model differs strongly, probably due to differences in the high-level concepts used by the vision models. Due to their lack of interpretability, the underlying mechanisms that cause the model to fail cannot be identified. The simple random forest models perform excellently in this generalization task. VGG11 and ResNet18 show high AUC ROC values compared to their relatively low balanced accuracy. This indicates that while these models are successful in differentiating the disease level, the classification threshold does not generalize to the new dataset.
V Interpretability Analysis

### v2

@@ 3.3.1 Attribution Method
Instead of using a constant baseline sample, such as a zero-vector, we propose a dynamic baseline [sturmfels2020visualizing] that considers the unique characteristics of retinal vasculature, generating more fine-grained explanations. In contrast to natural images, the retinal vasculature adheres to a fixed structure that can be represented in the explanation process. Previous studies indicate variations in capillary density with increasing distance from the FAZ and across different retinal sectors [lavia2020retinal]. This implies that distinct characteristics should be considered normal for intercapillary areas and vessels in different areas of the retinal vasculature. Therefore, we select the k-nearest same-type neighbor nodes from the training set and use their feature-wise mean as the dynamic baseline vector $x^{bl}$ for each node in a graph. Formally, the baseline vector $x_{v}^{bl}$ for a node $v$ is defined as
$x_{v}^{bl}=\text{{avg}}\left(\left\{x_{u}\,\middle|\,u\in\mathcal{N}_{l}(v)\right\}\right).$ (5)
@@ 3.3.2 Localization Method
For clinical utility, it is critical that graph nodes that have been identified as important (e.g., individual vessel segments) can be visualized in the corresponding OCTA image. These visualizations can be used to verify whether the extracted retinal characteristics align with the underlying image or if they arise from representation errors, such as image artifacts or segmentation errors. We store the corresponding segmentation map for each node in our graph representation to achieve such a localization.
3.3.3 Explanation Generation
@@
To generate our explanations, we first identify critical nodes using our attribution method by summing over the integrated gradients for the feature vector. The identified critical nodes are then spatially highlighted using the described localization method. Lastly, we identify the decisive characteristics of each critical node using the feature-wise integrated gradients. Examples of the generated spatial and feature-wise explanations are described in Section 4 and visualized in Figure 8.
3.4 Datasets
3.4.1 Proprietary Clinical Dataset
@@
Our OCTA dataset contains 1268 OCTA images of the DVC. For some patients, images of both the right and left eyes are contained in the dataset. Beyond that, no patients are included more than once. DR grades are assigned on a patient level with the stages of healthy (847 samples), diabetes mellitus (DM) without DR (133 samples), early non-proliferative DR (early NPDR, 134 samples), late NPDR (62 samples), and PDR (92 samples). The disease grades were assigned according to the International Clinical Disease Severity Scale [48], relying on color fundus photography. The OCTA images are 3 mm $\times$ 3 mm, with a resolution of $\times304$ pixels.
@@
We perform a stratified data split into six parts for our experiments. As sometimes both eyes of a single patient are imaged, we strictly keep both samples in the same split to avoid information leakage. One of the splits is kept as a test set, while the other five splits are used for cross-validation. In each split, healthy and DM are pooled into a single healthy class, and early NPDR and late NPDR are pooled into a single NPDR class. As is typical in a real-world setting, our datasets are imbalanced towards healthy subjects. For these reasons, we evaluate and select our models based on balanced agreement.
3.4.2 Public Dataset
We use the OCTA-500 [li2020octa] dataset as an external validation dataset for our method. For that purpose, we use the 200 3 mm $\times$ 3 mm images with a resolution of $\times304$ pixels and remove samples with labels other than DR or healthy. Labels in the OCTA-500 dataset are assigned based on medical patient records. This results in a total of 189 images, of which we use the ILM-OPL projection maps. Notably, these projections contain SVC and DVC vessels, resulting in a slight domain shift compared to our proprietary dataset. Of the 189 images, 160 are healthy samples, and 29 are in the DR class. We apply each algorithmic class’s best-performing checkpoints to classify the OCTA-500 samples. Predictions of NPDR and PDR are pooled into a single DR class for comparison with the ground truth labels.
3.5 Baseline Methods
We compare our approach to the most prominent and well-performing approaches for DR staging in the literature. These methods include ML using pre-established biomarkers and DL approaches such as CNNs directly operating on the information-rich raw OCTA images. Furthermore, we compare our method to an ophthalmology foundation model that has not previously been employed for this task on OCTA images. Finally, we establish traditional ML methods on graphs that employ global feature aggregation followed by classic ML classifiers without using graph convolutions.
3.5.1 Manual Biomarker Extraction
We extract traditional biomarkers from the image segmentations. We use RF and SVM as classifiers for the DR stage prediction due to their excellent performance in a wide range of applications for tabular data. Specifically, we extract the following biomarkers and use them as classification features:
  1. 1.
    FAZ biomarkers: area, maximal diameter, mean diameter, acircularity index, STD4, and NR300. The reader is referred to [lu2018evaluation] for detailed descriptions of the last three biomarkers.
  2. 2.
    Vessel and intercapillary area biomarkers: vessel density, vessel perimeter, and fractal dimension as a vessel complexity measure.
3.5.2 Vision Models
We compare to well-established CNN architectures previously used for the task at hand [qian2023drac]. We implemented EfficientNet [tan2019efficientnet], ResNet [xie2017aggregated], and VGG [simonyan2014very] models in combination with an optimized data augmentation pipeline. We use all models with weights pre-trained on ImageNet. Beyond these networks, we use the more recent ConvNeXt [liu2022convnet], Efficient Net V2 [tan2021efficientnetv2], and SwinTransfomer [liu2021swin] architectures. We use cross-entropy loss for CNNs and all other neural network architectures.
3.5.3 RETFound Foundational Model
The RETFound model [zhou2023foundation] is built on a vision transformer architecture. To initialize this complex architecture’s weights, the model is pre-trained in a self-supervised fashion on more than 900,000 color fundus photography images before being fine-tuned on our training set. We decided to use the color fundus pre-trained model instead of the OCT fine-tuned models due to better performance. This is likely caused by the larger domain shift from en-face OCTA to OCT than to color fundus photography.
3.5.4 Traditional ML on Node Embeddings
3.6 Metrics
We report ROC-AUC for no DR, NPDR, and PDR class staging and for the binary DR vs no DR classification. The class balanced agreement metric indicates how often the clinical fundus labeling and the OCTA-based stage prediction match. F1 scores are used to indicate the class-specific agreement rates. Perfect agreement rates for clinical fundus labels and 3mm x 3mm microvascular information are neither expected nor desired. This would disregard the additional microvascular information from OCTA imaging, see 5 and section 4.5.
@@ 3.6 Metrics
Table 2: Agreement scores for the staging task into fundus labels healthy, NPDR and PDR. The data representation describes the information used as input to the subsequent classifier. Results are obtained on a test set that was separated before cross-validation. Standard deviations are calculated based on the results from the 5-fold cross-validation. We report p-values for paired t-tests with the heterogeneous graph model in parentheses after the standard deviations.
@@
DR vs no DR
ROC AUC
Trad. ML
Tabular
Biomarkers
RF
@@
Transformer Image RETFound
@@
Trad. ML
Node
Embeddings
RF
@@
4 DR Staging Experiments
4.1 Agreement Rates of Machine Learning OCTA Grading Approaches
Our biology-informed heterogeneous graph representation achieves comparable ROC AUC scores of 0.829, matching the best CNN (0.822) and transformer models (0.819). Staging purely based on established biomarkers reaches notably lower agreement with the CFP-based labels. We attribute the lower agreement to the lack of biomarkers that fully encode pathological information on intraretinal microvascular abnormalities and microaneurysms. In contrast, CNNs and our heterogeneous graph representation plateau at agreement rates where the DR stage is determined by microvascular information in OCTA images that diverges from clinical evaluation through CFP (see Figure 5).
Table 3: Classification performance for the binary classification into healthy versus DR. The results are obtained on the external OCTA-500 without further fine-tuning the models previously trained on the other reported dataset. The checkpoint from the 5-fold cross-validation on the proprietary dataset with the highest balanced agreement is used for evaluation on the OCTA-500 dataset.
@@ 4.1 Agreement Rates of Machine Learning OCTA Grading Approaches
Trad. ML
Tabular
Biom.
RF 0.821 0.963 0.760 0.960 0.872
@@
4.2 Sensitivity analysis on OCTA image variations and segmentation maps
@@
A comparison of using the SVC instead of the DVC using the same CNN architecture shows similar staging performance for using either the DVC or the SVC layer. Moreover, we experimented with directly using segmentation maps as input to the CNN classifier. Interestingly, this results in comparable performance to raw images and opens an interesting avenue for enhancing explainability in classification with CNN models.
4.3 Generalization to public OCTA-500 dataset
@@
We use the checkpoints with the highest class-balanced agreement rates from the 5-fold cross-validation on the proprietary dataset for evaluation on the OCTA-500 dataset. Our proposed graph learning method yields the highest agreement with the medical record labels in this 0-shot evaluation on the OCTA-500 dataset (see Table 3). This highlights the robustness of our approach to minor domain changes induced by different hardware (OCTA scanners) and preprocessing algorithms (e.g., layer segmentation and intensity clipping). The performance between the CNN models and the RETFound model differs strongly, probably due to differences in the high-level concepts used by the vision models. Due to their lack of interpretability, the underlying mechanisms that cause the model to fail cannot be identified. The simple random forest models perform excellently in this generalization task. VGG11 and ResNet18 exhibit high AUC ROC values compared to their relatively low balanced agreement. This indicates that while these models are successful in differentiating the disease level, the classification threshold does not generalize to the new dataset.
Table 4: Comparison for using vessel graph, intercapillary area graph, and heterogeneous graph
Data
Representation
Model
ROC
AUC
Bal.
Agr.
F1
Healthy
F1
NPDR
F1
PDR
DR vs no DR
ROC AUC
Vessel Graph GNN
0.828
$\pm$0.005
0.642
$\pm$0.025
0.812
$\pm$0.037
0.372
$\pm$0.060
0.516
$\pm$0.048
0.876
$\pm$0.010
Intercapillary
Area Graph
GNN
0.815
$\pm$0.018
0.627
$\pm$0.043
0.836
$\pm$0.047
0.336
$\pm$0.178
0.498
$\pm$0.068
0.881
$\pm$0.012
Heterogeneous Graph GNN
0.829
$\pm$0.037
0.696
$\pm$0.034
0.819
$\pm$0.021
0.419
$\pm$0.057
0.568
$\pm$0.090
0.867
$\pm$0.003
4.4 Comparison to Homogeneous Graph Representations
We experimented with a simplified setting of the individual homogeneous graph representations and included the results below in Table 4. The intercapillary region graph does not achieve comparable agreement rates measured by ROC AUC (0.815), compared to both vessel (0.828) and heterogeneous graphs (0.829). We attribute this to the lack of information about intraretinal microvascular abnormalities and microaneurysms. The performance of the vessel graph is comparable to the performance of the heterogeneous graph. In fact, the vessel graph can indirectly encode the crucial pathological non-perfusion information through a reduction in total vessel area. However, only the heterogeneous graph allows for a direct attribution to the non-perfusion regions and therefore the absence of capillaries.
Figure 5: Overview of cases where the microvascular changes assessed using en face OCTA images do not align with the disease progression information obtained from color fundus photography (CFP). (top row) Comparison of an en face OCTA image with the matching CFP image (middle and bottom row). Further cases of OCTA images where the OCTA-based expert assessment does not align with CFP grading.
4.5 Expert reading of OCTA predictions that disagree with clinical fundus stage labels
DR stage labels in the proprietary dataset were assigned according to the International Clinical Disease Severity Scale [wilkinson2003proposed] based on the Early Treatment Diabetic Retinopathy Study (ETDRS), which relies on color fundus photography. This feature-based grading approach does not necessarily align with the microvasculature changes visible in OCTA images [thompson2019optical]. Pathological microvascular alterations can occur without microaneurysms observable on fundus photography, a phenomenon we encountered frequently in our dataset.
To investigate these staging discrepancies, we performed staging experiments with a senior ophthalmologist with over ten years of OCTA and DR experience. Nine cases were preselected where OCTA model predictions plausibly disagreed with fundus labels (see Section 5). Expert review confirmed the disagreement in eight of these cases. We report cases where expert ophthalmologists assess OCTA images as NPDR or even PDR stage, while no retinopathy is detected in the corresponding fundus images (see Figure 5 first and last row). Conversely, sometimes no microvascular abnormalities are visible in the spatially limited 3 mm x 3 mm OCTA image, while the color fundus examination indicates the DR (see Figure 5 middle row).
In clinical practice, such modality discordance between undus and OCTA labels is particularly relevant, as clinicians must reconcile information from both imaging techniques to formulate unified staging and treatment decisions.
5 Interpretability Analysis and Explainability of Staging
5.1 Feature Relation to Microvascular Pathologies
Our feature embeddings for intercapillary areas, vessel segments, and the FAZ are directly linked to pathological changes of the microvasculature. Figure 6 illustrates the relationship between specific features and the progression of DR. The increase in tortuosity of vessels is defined as a microvascular abnormality; we observe a clear increase in the 90% quantile tortuosity of vessel segments (Figure 6 top left). Capillary degeneration results in less branching and therefore larger vessel segments; moreover, microaneurysms and neovascularizations often appear as large segmented vessel segments. An increased 90% quantile for vessel segment volume clearly reflects these pathological changes (Figure 6 top center). Microvascular abnormalities and microaneurysms are characterized by irregular shapes. We observe that the 90% quantile of the radius variability of isolated vessel segments is increased for the NPDR and PDR stage (Figure 6 top right). Capillary loss is best reflected in the characteristics of intercapillary areas. We see a clear increase in the median area of intercapillary areas (Figure 6 middle left), and even more drastically for the 90% quantile (Figure 6 middle center). This capillary loss is also reflected in other characteristics, such as the increased median major axis length of intercapillary areas (Figure 6 middle right). Changes to the FAZ characteristics, due to the loss of integrity of parafoveal vessels, are a known indicator for the progression of DR. We see this effect most pronounced in the decreased solidity (defined as the ratio of the area of a region to the area of its convex hull) (Figure 6 bottom left). Changes, although less pronounced, are also visible in the area and the length of the minor axis of the FAZ (Figure 6 bottom center and right).
Figure 6: Analysis of specific features that are encoded in our heterogeneous graph representation for vessel segments (top row), intercapillary areas (middle row), and the FAZ (bottom row). Violin plots are stratified by clinical DR grade. The violin plots show the distribution of a specific characteristic (e.g., median intercapillary area size) across all OCTA scans of the specified disease stage.